Date:
Thursday, May 27, 2021
Telescope:
CTIO 1.5-m
Submitter:
Rodrigo Hinojosa
Submitter Email:
Observer Support:
None
Night Assistant:
Rodrigo Hinojosa
Problems:
None.
Program:
Target of Oportunity:
No
Time Observed:
0.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
11.5
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
11.5
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2021-05-27 | 0.0 | 0.0 | 0.0 | 11.5 | 0.0 | 11.5 |
Overcast and some precipitation last night. No observations done. Telescope is covered and dome facing north.
(Leonardo Paredes comments follow)
With Rodrigo we conducted test to recalculate the exposure times for quartz and ThAr lamps for the replaced fiber cable for the lamp box. It took us around 4 hours to setup an efficent and consistent measurement method of the order profiles (thanks all mighty iraf!), along with the subsequent exposure times trials.
- We found a consistent 45% decreased flux in quartz lamps compared to frames taken before the replacement, we compensated to match the previous flux by increasing the exposure times on the exp_times file definition:
- quartz 3x1 narrow_slit now increased from 40s to 58s
- quartz 3x1 slit now increased from 19s to 28s
- quartz 4x4 fiber now increased from 0.8s to 1.16s
- quartz 3x1 slicer now increased from 14s to 20s. Here we found a problem to match the previous flux. Only 2 out of the 3 sliced slits matched when compensating the exptime. The profile of the third slit had more missing flux.
The latter led us to suspect that the decreased flux could be due to a non-uniform illumination of the fiber. Could be some misalignment or maybe the fiber has a different specification compared to the previous one?
We also conducted test on the ThAr exposures. Compared to previous nights, the continuum (or maybe bias level) was higher and the lines were shallower, making harder to measure the amount of flux missing. Which explained also the lower number of ThAr lines found useful for fwhm when focusing the instrument few days ago. We increased the exposure times on the ThAr lamp to match the peak of the previous lines, despite the continuum level being increased. We settled with an 80% increase.
- ThAr 3x1 narrow_slit increased from 0.8s to 1.44s
- ThAr 3x1 slit increased from 0.4s to 0.72s
- ThAr 3x1 fiber increased from 0.19s to 0.342s
- ThAr 3x1 slicer increased from 0.4s to 0.72s. For now, hard to tell if the same inconsistency between the slices was present.
At this 80% increased exposure time, the number of ThAr lines found useful for fwhm measures in the focus routine now matched closely the number found before the fiber replacement (~190 for ThAr narrow_slit).
The end of night calibrations we left running with these new exposure times. Automatic afternoon calibrations were disabled so we can work with Andrei and analyse the data.